OpenStudy (death47):

Which fraction is equivalent to 3/7? A.15/35 B.6/10 C.9/14 D.21/42

OpenStudy (death47):

@Susanna @marihelenh

OpenStudy (marihelenh):

Start with A. Do you know the GCF?

OpenStudy (susanna):

times both sides by 5

OpenStudy (death47):

I think its A.

OpenStudy (marihelenh):

You would be correct!

OpenStudy (susanna):

yes

OpenStudy (mathstudent55):

No, divide both the numerator and denominator by 5 of choice A.

OpenStudy (death47):

OK, thanks guys. I think I'm getting it now :D (its been a while since ive done math...its usually English related what I do ;p)

OpenStudy (mathstudent55):

You need to reduce all the fractions and then you can compare them with the given fraction.

OpenStudy (mathstudent55):

A. What is the largest number you can divide both 15 and 35 evenly by?

OpenStudy (death47):

5?

OpenStudy (susanna):

yes

OpenStudy (mathstudent55):

Correct. Divide both 15 and 35 by 5 to find its reduced fraction. What do you get?

OpenStudy (death47):

3/7?

OpenStudy (mathstudent55):

A. 15/35 = 3/7 Correct.

OpenStudy (mathstudent55):

Now for practice and completeness, let's try B. B. What is the largest number you can divide both 6 and 10 evenly by?

OpenStudy (death47):

I think the highest number they have in common is 2, right?

OpenStudy (mathstudent55):

Correct. B. 6/10 = 3/5

OpenStudy (mathstudent55):

We already know A is the answer, since 15/35 = 3/7. B is 3/5 which is not 3/7. If you want, we can continue with C and D so you get a little more practice reducing fractions.

OpenStudy (mathstudent55):

C.9/14 What is the largest number that 9 and 14 are divisible by?

OpenStudy (death47):

yea, let's, id like to get some more practice before I continue :) better safe than sorry.

OpenStudy (mathstudent55):

C.9/14 What is the largest number that 9 and 14 are divisible by?

OpenStudy (mathstudent55):

9 = 3 * 3 14 = 2 * 7 As you can see, 9 and 14 have no prime factors in common. That means the largest number 9 and 14 can be divided by evenly is 1. That means that 9/14 cannot be reduced since it is already a reduced fraction.

OpenStudy (death47):

I thought 3...but 3 cant go into 14...so 1? but everything is divisible by 1, sooo.....^^lol

OpenStudy (mathstudent55):

Right. Now we can move onto D. 21/42

OpenStudy (death47):

that would be 21, right?

OpenStudy (mathstudent55):

Exactly.

OpenStudy (mathstudent55):

Divide both 21 and 42 by 21. What do you get?

OpenStudy (death47):

1/2

OpenStudy (mathstudent55):

Correct. Now we can review the entire problem. Which fraction is equivalent to 3/7? A. 15/35 = 3/7 B. 6/10 = 3/5 C. 9/14 D. 21/42 = 1/2 With all fractions reduced, we see the answer is A. 15/35.

OpenStudy (mathstudent55):

Great job!

OpenStudy (death47):

Thank you so much :D

OpenStudy (mathstudent55):

yw